3. In December, RMO Manufacturing had Actual Output of 100 units. Its Effective Capacity dropped to 120 units and its Design Capacity remained constant at 200 units. Calculate the Efficiency and Utilization for December. Compare the December rates to the September rates? Is RMO more or less efficient? Is this a good or a bad thing? Explain your answer
Answer #1

DESIGN CAPACITY = 200
EFFECTIVE CAPACITY = 120
ACTUAL CAPACITY = 100

UTILIZATION = ACTUAL CAPACITY / DESIGN CAPACITY = 100 / 200 = 0.5 OR 50%
EFFICIENCY = ACTUAL CAPACITY / EFFECTIVE CAPACITY = 100 / 120 = 0.8333 OR 83.33%


EVEN THOUGH THE CAPACITY UTILIZATION HAS DROPPED QUITE A BIT, THE EFFICIENCY IS APPROPRIATE AT 83.3%

EFFICIENCY IS A FACTOR OF THE USE OF CAPACITY BASED ON THE EFFECTIVE CAPACITY, WHICH COULD BE LOWER THAT THE DESIGN CAPACITY DUE TO SERVICE, REPAIRS, PRESENCE OF VARIOUS PRODUCTS, ETC.
WHILE UTILIZATION IS BASED ON THE DESIGN CAPACITY, EFFICIENCY IS AT APPROPRIATE LEVELS

** Leaving a thumbs-up would really help me out. Let me know if you face any problems.